OpenStudy (anonymous):

f(x) = log2[(x+1)/(x-1)] dins the derivative

OpenStudy (australopithecus):

find derivative?

OpenStudy (anonymous):

yes

OpenStudy (australopithecus):

I would split the logarithm f(x) = log_2(x+1) - log_2(x-1) then just use the basic rule log_n(x) = x'/xlog(n)

OpenStudy (australopithecus):

you can use chain rule if you want

OpenStudy (anonymous):

show me how to solve it using chain rule

OpenStudy (australopithecus):

so you would have f'(x) = 1/(x+1)log(2) - 1/(x-1)log(2)

OpenStudy (australopithecus):

ok I can show you how to use chain rule but give me some time

OpenStudy (anonymous):

ok sure :)

OpenStudy (australopithecus):

Do you want me to use product rule or quotient rule

OpenStudy (australopithecus):

let me know now or I'm just going to pick :L

OpenStudy (anonymous):

use tthe quotient rule

OpenStudy (anonymous):

anything no problem

OpenStudy (australopithecus):

f(x) = log2[(x+1)/(x-1)] So to deal with chain rule (and most other rules it can help as well such as product etc..) the best course of action when you are new to derivatives is to split the function into smaller functions so set l(x) to the outer function and g(x) to the inner function l(x) = log_2(x) l'(x) = 1/xlog(2) g(x) = (x+1)/(x-1) We need to use quotient rule or product rule to find g'(x) so split it up into smaller functions s(x) = (x+1) s'(x) = 1 d(x) = (x-1) d'(x) = 1 so now I'm going to apply quotient rule to solve g'(x) which is, ( s'(x)d(x) - s(x)d'(x) )/(d(x))^2 so, g'(x) = ( s'(x)d(x) - s(x)d'(x) )/(d(x))^2 g'(x) = ( (x-1) - (x+1) )/(x-1)^2 so now we have, l(x) = log_2(x) l'(x) = 1/xlog(2) g(x) = (x+1)/(x-1) g'(x) = ( (x-1) - (x+1) )/(x-1)^2 sub that into chain rule to solve for f'(x) chain rule being l'(g(x))*g'(x) so we have f'(x) = l'(g(x))*g'(x) f'(x) = ( 1/((x+1)/(x-1))log(2) ) * ( (x-1) - (x+1) )/(x-1)^2

OpenStudy (australopithecus):

eventually you should be able to do this in your head but it can still help to write it out to make sure you do not make mistakes when derivatives get really dirty

OpenStudy (anonymous):

thak you :))

OpenStudy (australopithecus):

I don't understand why most teachers don't teach derivatives this way